# Hybrid collective excitations in topological superconductor/ferromagnetic insulator heterostructures

### v3

@@ IV.1 Higgs mode
In Appendix B it is shown that $A_{12}^{K*}(-\omega^{*},-\bm{k})=A_{12}^{K}(\omega,\bm{k})$, $B_{12}^{K*}(-\omega^{*},-\bm{k})=B_{12}^{K}(\omega,\bm{k})$ and $C_{12}^{K*}(-\omega^{*},-\bm{k})=C_{12}^{K}(\omega,\bm{k})$. Making use of these symmetry relations from Eq. (62) we immediately obtain that $F_{\omega,\bm{k}}^{a}=0$ for the TS. It means that if the TS is in the helical ground state the amplitude Higgs mode is not coupled to the magnon in the linear order.
@@
Figure 2: Real and imaginary parts of the Higgs mode susceptibility $\chi_{H}(\omega,k)$ as a function of the excitation frequency at $Dk^{2}/\Delta=0.25$ and $T=0.1T_{c}$ (orange lines) in comparison to the corresponding results taken from Ref. [24] (black points).
@@ IV.2 NG mode in the absence of the coupling to FI
Figure 3: The spectrum (a) and decay rate (b) of superconducting phase mode at various plasma frequencies calculated at $\Gamma=0.018\Delta$ and $T=0.1T_{c}$, which is shown as a horizontal dashed line. Color dashed lines in panel (a) are results of analytical calculation making use of Eq. (66).
The dispersion $\omega(k)$ and the decay rate $\kappa(k)$ of the NG mode calculated according to this equation are shown in Fig. 3 for different plasma frequencies. At low temperatures $T\ll\Delta$, low frequencies $\omega\ll\Delta$ and small wavenumbers $k\xi\ll 1$ the spectrum of the NG mode of the charged 2D superconductor can be found analytically. For this purpose we follow the procedure of Ref.[85] and arrive at the following matrix equation
$\begin{pmatrix}\frac{(\hbar\omega)^{2}-\pi\Delta_{0}\hbar Dk^{2}}{4\Delta_{0}^{2}}&\frac{i\hbar\omega}{2\Delta_{0}}\\ -\frac{i\hbar\omega}{2\Delta_{0}}&1+\frac{k}{2\pi e^{2}\nu}\end{pmatrix}\begin{pmatrix}\delta\Delta_{p}\\ e\delta\phi_{\omega,k}\end{pmatrix}=0.$~~~~ (65)
@@
The analytical expression Eq. (66) is in reasonable agreement with our numerical result presented in Fig. 3(a) in the appropriate range of parameters $\omega\ll\Delta$ and $k\xi\ll 1$. The dispersion relation of the NG mode demonstrates a crossover between the linear and square root in momentum behavior with increase of the superconducting plasma frequency.
@@ IV.3 Hybridized magnon-NG mode
Figure 4: Hybridization between NG and magnon modes. (a) - the excitation propagates along the $x$-axis and (b)- along the $y$-axis. $\delta\omega_{a}$ is the anticrossing strength, which is defined in Eq. (IV.3).
@@
$\begin{pmatrix}M_{\omega,\bm{k}}^{p}&{\rm sgn}k_{x}C_{12}^{K}(\omega,\bm{k})\\ c_{s}{\rm sgn}(k_{x})J_{p}&M_{\omega,\bm{k}}^{h}\end{pmatrix}\hat{\Psi}_{p}=0,$ (67)
where $M_{\omega,\bm{k}}^{p}=i\left[A_{12}^{K}(\omega,\bm{k})-1-B_{12}^{K}(\omega,\bm{k})\right]$, $J_{p}=i\left[J_{\delta\Delta}(\omega,\bm{k})-J_{\delta\Delta^{*}}(\omega,\bm{k})\right]$ and $M_{\omega,\bm{k}}^{h}=\omega^{2}/(\omega_{b}+i\alpha\omega)-(\omega_{b}+i\alpha\omega)+c_{s}[J_{0,\omega}+{\rm sgn}(k_{x})J_{2,(\omega,\bm{k})}]$.
The dispersion curves of the hybridized magnon-NG mode were calculated numerically from Eq. (67) in the Companion paper [87] and are shown here in Fig. 4 just for completeness and the benefit of the reader. First of all, due to the spin-momentum locking which dictates the symmetry of the linear response of the OP on the magnon in the form $(\hat{\bm{k}}\times\delta\bm{h}_{\omega,\bm{k}})\cdot\hat{z}$ the coupling is anisotropic. Its magnitude is maximal if the magnon propagates along the equilibrium magnetization direction ($x$-axis), leading to the anticrossing between the NG and the magnon modes, and is zero if the magnon propagates along the $y$-axis, as it is demonstrated in Fig. 4. Here we focus on the analytical description of spectrum in the vicinity of the anticrossing region for the case of maximal hybridization $\hat{\bm{k}}=\hat{x}$.
@@
Assuming magnetic parameters corresponding to YIG $\gamma=1.76\cdot 10^{7}G^{-1}s^{-1}$, $\omega_{0}=\gamma K=10^{-17}erg\approx 0.018\Delta_{0}$, $M_{S}=140G$, and $D_{m}=5\cdot 10^{-29}erg\cdot cm^{2}$ [89], we obtain that the hybridization region corresponds to the limit $k_{i}\xi\ll 1$, as it is also seen in Fig. 4.
@@
Figure 5: Anticrossing strength $\delta\omega_{a}=\omega_{up}(k_{i})-\omega_{dn}(k_{i})$ as a function of $\omega_{0}$ (a), $\omega_{p}$ (b), effective exchange field $h_{0}$ (c) and temperature $T$ (d). In plots (a) and (c) $\omega_{p}=4.35\Delta$, while $h_{0}\approx 1.17\Delta$ in plots (a), (b) and (d).
@@
The anticrossing strength $\delta\omega_{a}$ calculated numerically from Eq. (67) without making use of the approximations of low temperature, low frequencies and small wavenumbers, is shown in Fig. 5 as function of all essential physical parameters. The numerical results support our analytical low-temperature findings. As it is seen from Figs. 5(a)-(b) at low temperatures $\delta\omega_{a}$ does not depend on the magnon frequency $\omega_{0}$ and superconducting plasma frequency $\omega_{p}$, although this statement is violated at higher temeperatures. At the same time the linear dependence of $\delta\omega_{a}$ on $h_{0}$ survives in the whole temperature range. It is the most general result, which follows directly from Eq. (67) since $\delta\omega_{a}\propto\sqrt{M_{\omega,\bm{k}}^{p}F_{\omega,\bm{k}}^{p}c_{s}(\bm{J}_{p}\cdot\hat{x})}\propto h_{0}$ because the other involved quantities do not contain $h_{0}$.
@@ V Conclusion
A linear response theory of collective excitations in disordered 2D superconducting systems having the property of the full spin-momentum locking contacted with a thin-film ferromagnetic insulator is developed. The theory is based on the nonequilibrium Keldysh-Usadel quasiclassical approach. Making use of the developed approach it is predicted that the magnons in the FI and the Nambu-Goldstone (NG) phase mode in the TS are coupled forming composite magnon-NG excitation. The hybridization occurs via the interface exchange coupling between the conductivity electrons of the 2D TS superconducting surface state and the FI magnetization. Due to the spin-momentum locking of electrons in the helical surface state of the TS the superconducting condensate has the same magnitude of singlet and triplet correlations, thus giving the superconducting OP the ability to respond to a magnon. On the other hand, excitation of a phase mode in the TS leads to the appearance of ac current, which is always accompanied by electron spin polarization (direct magnetoelectric effect). The current-induced spin polarization creates a torque, generating magnons in the FI. The coupling strength is studied analytically and numerically. It is demonstrated that the coupling strength depends linearly on the interface exchange coupling constant and also depends on the magnon gap exhibiting square root behavior at low temperatures. The dependence on the superconducting plasma frequency is not essential at low temperatures or at high values of the plasma frequency. The coupling of the magnon to the amplitude Higgs mode is also studied. It is demonstrated that the Higgs mode is not coupled to magnons in the linear order and does not form hybrid collective excitations with it.
@@ Appendix A First-order corrections to the Green’s functions
$[\delta\hat{g}_{\omega,k}^{R,A}]_{ij}=a_{ij}^{R,A}\delta\Delta_{\omega,k}+b_{ij}^{R,A}\delta\Delta_{-\omega,-k}^{*}$
$+c_{ij}^{R,A}(\hat{\bm{k}}\times\delta\bm{h})\cdot\hat{z}+d_{ij}^{R,A}\delta\phi,$ (80)
@@
$\left[\delta g_{\omega,k}^{K}\right]_{11}=\frac{1}{D_{n}^{K}}\left(-a_{2}b_{2}+a_{3}b_{3}+a_{4}b_{1}\right),$ (93)
$\left[\delta f_{\omega,k}^{K}\right]_{12}=\frac{1}{D_{n}^{K}}\left(a_{1}b_{2}-a_{2}b_{1}+a_{3}b_{4}\right),$ (94)
$\left[\delta f_{\omega,k}^{K}\right]_{21}=\frac{1}{D_{n}^{K}}\left(-a_{1}b_{3}-a_{2}b_{4}+a_{3}b_{1}\right),$ (95)
$\left[\delta g_{\omega,k}^{K}\right]_{22}=\frac{1}{D_{n}^{K}}\left(-a_{2}b_{3}+a_{3}b_{2}-a_{4}b_{4}\right),$ (96)
@@
$b_{1}=$ $-\delta\Delta_{\omega,k}f_{0}^{K}(\epsilon+\frac{\omega}{2})+\delta\Delta^{*}_{-\omega,-k}f_{0}^{K}(\epsilon-\frac{\omega}{2})-i\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xx}^{K}-g_{-}^{K}\delta\phi_{\omega,k}$
@@
$b_{2}=$ $-i\delta\Delta_{\omega,k}g_{+}^{K}+\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xy}^{K}+2i\Gamma\left(\tanh{\frac{\epsilon+\omega/2}{2T}}\delta{f}^{R}_{12}+\tanh{\frac{\epsilon-\omega/2}{2T}}\delta{f}^{A}_{12}\right)-if_{-}^{K}\delta\phi_{\omega,k}$
@@
$b_{3}=$ $i\delta\Delta^{*}_{-\omega,-k}g_{+}^{K}+\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xy}^{K}-2i\Gamma\left(\tanh{\frac{\epsilon+\omega/2}{2T}}\delta{f}^{R}_{21}+\tanh{\frac{\epsilon-\omega/2}{2T}}\delta{f}^{A}_{21}\right)-if_{-}^{K}\delta\phi_{\omega,k}$
@@
$b_{4}=$ $\delta\Delta_{\omega,k}f_{0}^{K}(\epsilon-\frac{\omega}{2})-\delta\Delta^{*}_{-\omega,-k}f_{0}^{K}(\epsilon+\frac{\omega}{2})+i\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xx}^{K}+g_{-}^{K}\delta\phi_{\omega,k}$

### v4

@@ IV.1 Higgs mode
In Appendix B it is shown that $A_{12}^{K*}(-\omega^{*},-\bm{k})=A_{12}^{K}(\omega,\bm{k})$, $B_{12}^{K*}(-\omega^{*},-\bm{k})=B_{12}^{K}(\omega,\bm{k})$ and $C_{12}^{K*}(-\omega^{*},-\bm{k})=C_{12}^{K}(\omega,\bm{k})$. Making use of these symmetry relations from Eq. (68) we immediately obtain that $F_{\omega_{r},\bm{k}}^{a}=0$ for the TS. It means that if the TS is in the helical ground state the amplitude Higgs mode is not coupled to the magnon in the linear order.
@@
Figure 2: Real and imaginary parts of the Higgs mode susceptibility $\chi_{H}(\omega,k)$ as a function of the excitation frequency at $Dk^{2}/\Delta_{0}=0.25$ and $T=0.1T_{c}$ (orange lines) in comparison to the corresponding results taken from Ref. [65] (black points).
@@ IV.2 NG mode in the absence of the coupling to FI
Figure 3: The spectrum (a) and decay rate (b) of superconducting phase mode at various plasma frequencies calculated at $\Gamma=0.018\Delta_{0}$ and $T=0.1T_{c}$, which is shown as a horizontal dashed line. Color dashed lines in panel (a) are results of analytical calculation making use of Eq. (72).
Figure 4: The spectrum of superconducting phase mode at various temperatures: (a) - $\hbar\omega_{p}=1.79\Delta_{0}$ and $\Gamma=0.0018\Delta_{0}$, (b) - $\hbar\omega_{p}=1.79\Delta_{0}$ and $\Gamma=0.018\Delta_{0}$, (c) - $\hbar\omega_{p}=4.28\Delta_{0}$ and $\Gamma=0.0018\Delta_{0}$, (d) - $\hbar\omega_{p}=4.28\Delta_{0}$ and $\Gamma=0.018\Delta_{0}$.
The dispersion $\omega(k)$ and the decay rate $\kappa(k)$ of the NG mode calculated according to this equation are shown in Fig. 3 for different plasma frequencies. At low temperatures $T\ll\Delta$, low frequencies $\omega\ll\Delta$, small wavenumbers $k\xi\ll 1$ and $\Gamma\ll\omega$ the spectrum of the NG mode of the charged 2D superconductor can be found analytically. For this purpose we follow the procedure of Ref.[141] and arrive at the following matrix equation
$\begin{pmatrix}\frac{(\hbar\omega_{r})^{2}-\pi\Delta_{0}\hbar Dk^{2}}{4\Delta_{0}^{2}}&\frac{i\hbar\omega_{r}}{2\Delta_{0}}\\ -\frac{i\hbar\omega_{r}}{2\Delta_{0}}&1+\frac{k}{2\pi e^{2}\nu}\end{pmatrix}\begin{pmatrix}\delta\Delta_{p}\\ e\delta\phi_{\omega_{r},k}\end{pmatrix}=0.$~~~~ (71)
@@
The analytical expression Eq. (72) is in reasonable agreement with our numerical result presented in Fig. 3(a) in the appropriate range of parameters $\omega\ll\Delta$, $k\xi\ll 1$ and $\Gamma\ll\omega$. The dispersion relation of the NG mode demonstrates a crossover between the linear and square root in momentum behavior with increase of the superconducting plasma frequency. We note that the collective excitations of the phase in dirty superconductors were extensively studied in Ref.[98].
In Fig.4 we show results for the NG mode calculated at wide temperature range. It is seen that at low temperature and small $\Gamma$ this mode is close to the square-root behavior (Fig. 4(a) and (c), black curves), which indicates the NG-plasmon coupling. At the same time, at higher temperatures or at higher values of $\Gamma$ (see other curves in Fig. 4) the behavior gradually evolves from square-root to linear, which means that the NG mode becomes uncoupled from the plasmon. This is due to the presence of thermally-induced quasiparticles at high temperatures or even at low temperatures if the gap is smeared by substantial $\Gamma$. Then the supercurrents induced by the phase mode are counterbalanced by the normal currents from thermally excited quasiparticles, without any need to generate electron density oscillations. In this case the gapless phase excitation is called the Carlson-Goldman mode and has a linear dispersion law [88, 89, 91, 98].
@@ IV.3 Hybridized magnon-NG mode
Figure 5: Hybridization between NG and magnon modes. Left column: $\hbar\omega_{p}=4.28\Delta_{0}$; right column - $\hbar\omega_{p}=42.8\Delta_{0}$. Top row: excitations propagating along the $x$-axis; bottom row: along the $y$-axis. The anticrossing strength $\delta\omega_{a}$ is defined in Eq. (IV.3). The parameters used for the calculations are as follows. Superconducting subsystem: $\Gamma=0.018\Delta_{0}$, $T=0.1T_{c}$, $\xi=3.5nm$ [142] and $\sigma=3\cdot 10^{14}c^{-1}$ (normal state conductivity). Magnetic subsystem: $\gamma=1.76\cdot 10^{7}G^{-1}s^{-1}$, $\omega_{0}=\gamma K=10^{-17}erg\approx 0.018\Delta_{0}$, $D_{m}=5\cdot 10^{-29}erg\cdot cm^{2}$ [143] and $4\pi M_{s}\approx 2\cdot 10^{3}Oe$ [144]. The exchange field is set to $h_{0}\approx 4.16\Delta_{0}$. Insets in panels (a) and (b) display the anticrossing strength $\delta\omega_{a}$ as a function of the temperature.
@@
$\begin{pmatrix}M^{p}&{\rm sgn}k_{x}C_{12}^{K}(\omega,\bm{k})\\ c_{s}{\rm sgn}(k_{x})J_{p}&M^{h}\end{pmatrix}\hat{\Psi}_{p}=0,$ (73)
where $M^{p}=i\left[A_{12}^{K}(\omega,\bm{k})-1-B_{12}^{K}(\omega,\bm{k})\right]$, $J_{p}=i\left[J_{\delta\Delta}(\omega,\bm{k})-J_{\delta\Delta^{*}}(\omega_{,}\bm{k})\right]$ and $M^{h}=\omega^{2}/(\omega_{b}+i\alpha\omega)-(\omega_{b}+i\alpha\omega)+c_{s}[J_{0,\omega_{r}}+{\rm sgn}(k_{x})J_{2,(\omega_{r},\bm{k})}]$.
The dispersion curves of the hybridized magnon-NG mode were calculated numerically from Eq. (73) are shown here in Fig. 5. First of all, due to the spin-momentum locking which dictates the symmetry of the linear response of the OP on the magnon in the form $(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}$ the coupling is anisotropic. Its magnitude is maximal if the magnon propagates along the equilibrium magnetization direction ($x$-axis), leading to the anticrossing between the NG and the magnon modes, and is zero if the magnon propagates along the $y$-axis, as it is demonstrated in Fig. 5. Here we focus on the analytical description of spectrum in the vicinity of the anticrossing region for the case of maximal hybridization $\hat{\bm{k}}=\hat{x}$.
@@
Assuming magnetic parameters corresponding to YIG $\gamma=1.76\cdot 10^{7}G^{-1}s^{-1}$, $\omega_{0}=\gamma K=10^{-17}erg\approx 0.018\Delta_{0}$, $4\pi M_{s}\approx 2\cdot 10^{3}Oe$ [144] , and $D_{m}=5\cdot 10^{-29}erg\cdot cm^{2}$ [143], we obtain that the hybridization region corresponds to the limit $k_{i}\xi\ll 1$, as it is also seen in Fig. 5.
@@
Figure 6: Anticrossing strength $\delta\omega_{a}=\omega_{up}(k_{i})-\omega_{dn}(k_{i})$ as a function of $\omega_{0}$ (a), $\omega_{p}$ (b), effective exchange field $h_{0}$ (c) and temperature $T$ (d). In plots (a) and (c) $\omega_{p}=4.28\Delta_{0}$, while $h_{0}\approx 4.16\Delta_{0}$ in plots (a), (b) and (d). All other parameters are consistent with Fig.5.
@@
The anticrossing strength $\delta\omega_{a}$ calculated numerically from Eq. (73) without making use of the approximations of low temperature, low frequencies and small wavenumbers, is shown in Fig. 6 as function of all essential physical parameters. The numerical results support our analytical low-temperature findings. As it is seen from Figs. 6(a)-(b), at low temperatures $\delta\omega_{a}$ does not depend on the magnon frequency $\omega_{0}$ and superconducting plasma frequency $\omega_{p}$, although this statement is violated at higher temperatures. At the same time the linear dependence of $\delta\omega_{a}$ on $h_{0}$ survives in the whole temperature range. It is the most general result, which follows directly from Eq. (73) since $\delta\omega_{a}\propto\sqrt{M^{p}F_{\omega_{r},\bm{k}}^{p}c_{s}(\bm{J}_{p}\cdot\hat{x})}\propto h_{0}$ because the other involved quantities do not contain $h_{0}$.
@@ V Conclusion
We develop a linear response theory for collective excitations in a disordered 2D superconductor with full spin-momentum locking, interfaced with a thin-film ferromagnetic insulator (FI). Our approach is based on the nonequilibrium Keldysh-Usadel quasiclassical formalism, applied to the surface state of a topological superconductor (TS).
We predict that magnons in the TS and the Nambu-Goldstone phase mode in the TS can hybridize to form composite magnon – NG excitations. This coupling represents a previously unexplored form of dynamical proximity effect in S/F heterostructures. The interaction arises via interfacial exchange coupling between conduction electrons in the 2D superconducting surface state of the TS and the magnetization of the FI. A key ingredient enabling this hybridization is the spin–momentum locking inherent to the helical surface state of the TS, which enforces equal magnitudes of singlet and triplet pairing correlations. As a result, the superconducting order parameter becomes sensitive to the magnon field. Conversely, excitation of the NG mode in the TS generates an ac supercurrent, which, due to spin–momentum locking, is accompanied by spin polarization (direct magnetoelectric effect). This current-induced spin polarization exerts a torque on the FI, thereby exciting magnons.
The magnon-NG coupling strength is investigated both analytically and numerically. We find that it scales linearly with the interface exchange constant and exhibits a square-root dependence on the magnon gap at low temperatures. The coupling shows no significant dependence on the superconducting plasma frequency in the low-temperature or high-frequency regimes. In contrast, the Higgs (amplitude) mode shows no linear coupling to the magnons and does not participate in the formation of hybrid excitations.
The reconstructed spectra of these composite excitations offer a potential route to experimentally probe the NG mode and its coupling to magnons. Beyond its fundamental interest, the predicted effect holds promise for superconducting spintronics, as it provides a new mechanism for mutual conversion between spin signals and spinless collective superconducting excitations.
@@ Appendix A First-order corrections to the Green’s functions
$[\delta\hat{g}_{\omega_{r},k}^{R,A}]_{ij}=a_{ij}^{R,A}\delta\Delta_{\omega_{r},k}+b_{ij}^{R,A}\delta\Delta_{-\omega_{r},-k}^{*}$
$+c_{ij}^{R,A}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}+d_{ij}^{R,A}\delta\phi_{\omega_{r},\bm{k}},$ (86)
@@
$\left[\delta g_{\omega_{r},k}^{K}\right]_{11}=\frac{1}{D_{n}^{K}}\left(-a_{2}b_{2}+a_{3}b_{3}+a_{4}b_{1}\right),$ (99)
$\left[\delta f_{\omega_{r},k}^{K}\right]_{12}=\frac{1}{D_{n}^{K}}\left(a_{1}b_{2}-a_{2}b_{1}+a_{3}b_{4}\right),$ (100)
$\left[\delta f_{\omega_{r},k}^{K}\right]_{21}=\frac{1}{D_{n}^{K}}\left(-a_{1}b_{3}-a_{2}b_{4}+a_{3}b_{1}\right),$ (101)
$\left[\delta g_{\omega_{r},k}^{K}\right]_{22}=\frac{1}{D_{n}^{K}}\left(-a_{2}b_{3}+a_{3}b_{2}-a_{4}b_{4}\right),$ (102)
@@
$b_{1}=$ $-\delta\Delta_{\omega_{r},k}f_{0}^{K}(\epsilon+\frac{\omega}{2})+\delta\Delta^{*}_{-\omega_{r},-k}f_{0}^{K}(\epsilon-\frac{\omega}{2})-i\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xx}^{K}-g_{-}^{K}\delta\phi_{\omega_{r},k}$
@@
$b_{2}=$ $-i\delta\Delta_{\omega_{r},k}g_{+}^{K}+\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xy}^{K}+2i\Gamma\left(\tanh{\frac{\epsilon+\omega/2}{2T}}\delta{f}^{R}_{12}+\tanh{\frac{\epsilon-\omega/2}{2T}}\delta{f}^{A}_{12}\right)-if_{-}^{K}\delta\phi_{\omega_{r},k}$
@@
$b_{3}=$ $i\delta\Delta^{*}_{-\omega_{r},-k}g_{+}^{K}+\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xy}^{K}-2i\Gamma\left(\tanh{\frac{\epsilon+\omega/2}{2T}}\delta{f}^{R}_{21}+\tanh{\frac{\epsilon-\omega/2}{2T}}\delta{f}^{A}_{21}\right)-if_{-}^{K}\delta\phi_{\omega_{r},k}$
@@
$b_{4}=$ $\delta\Delta_{\omega_{r},k}f_{0}^{K}(\epsilon-\frac{\omega}{2})-\delta\Delta^{*}_{-\omega_{r},-k}f_{0}^{K}(\epsilon+\frac{\omega}{2})+i\frac{\delta h}{2v_{f}}Dk_{x}\Gamma_{xx}^{K}+g_{-}^{K}\delta\phi_{\omega_{r},k}$